OpenStudy (melaniechaim):

An object falls 9 meters during the first second, another 27 meters during the second second, and then 81 meters more during the third second. If this pattern continues, how far does the object fall during the first 4 seconds? Hint: cap s sub n equals start fraction a sub one left parenthesis one minus r to the power of n end power right parenthesis over one minus r end fraction comma r ≠ 1, where a1 is the first term and r is the common ratio

OpenStudy (sshayer):

a=9 m r=27/9=3 \[S _{n}=a \frac{ r^n-1 }{ r-1 }\] put n=4 s and find S4

OpenStudy (melaniechaim):

I really just need the answer bc this is due in like 5 minutes